A local drugstore owner knows that, on average, 100 people enter his store each hour.


a) Find the probability that in a given 3-minute period nobody enters the store.


b) Find the probability that in a given 3-minute period more than 5 people enter the


store.

1
Expert's answer
2022-03-27T16:05:21-0400

We have a Poisson distribution,

"\lambda=100;
P_t(X=k)=\cfrac{(\lambda t)^k\cdot e^{-\lambda t}}{k!};\\
3\ \rm{mins=\cfrac{3}{60} \ hour=0.05\ hour;}"

"\rm For\ a\ 3-minute\ period \ \lambda t=100\cdot0.05=5."


a)

"P_{0.05}(X=0)=\cfrac{5^0\cdot e^{-5}}{0!}=0.0067."


b)

"P_{0.05}(X>5)=1-P_{0.05}(X\leq5)=\\
=1-(P_{0.05}(X=0)+P_{0.05}(X=1)+\\
+P_{0.05}(X=2)+P_{0.05}(X=3)+\\
+P_{0.05}(X=4)+P_{0.05}(X=5))=\\
=1-\\
-\begin{pmatrix}
 \cfrac{5^0}{0!}+\cfrac{5^1}{1!}+\cfrac{5^2}{2!}+\cfrac{5^3}{3!}+\cfrac{5^4}{4!}+\cfrac{5^5}{5!}\end{pmatrix}\cdot e^{-5}=\\
=1-\\
-(1+5+12.5+20.83+26.04+26.04)\cdot e^{-5}=\\
=1-0.6160=0.3840."